OpenStudy (anonymous):

Which of the following is the best step to make first when simplifying the expression ? 0.6-1/3

OpenStudy (solomonzelman):

\[0.6-1/3\]\[0.6=6/10=3/5\]\[1/3+1/5\]\[5/15+3/15\]\[8/15\]